Let f(x)=x^3 -2. What is f(1/2)?
Answer by Prithwis(166)

Find f(1/2) for f(x)=x^3 -2
f(1/2)=(1/2)^3 -2
=> f(1/2)=(1/8) -2
[Note: For (1/2)^3 just raise top and bottom to power of 3 and get 1/8]
=>f(1/2)= (1-16)/8
=>f(1/2)= (-15)/8
So, f(1/2)=(-15)/8 for when f(x)=x^3 -2